OpenStudy (anonymous):

In how many ways can seven identical candies distributed to 3 children {I, II, III} if each child can get any number of candies?

OpenStudy (amistre64):

isnt that just a count in base 7? 000 001 002 ... 007 010 011 012 .... or do all the candies have to be handed out at once?

OpenStudy (amistre64):

i spose it would have to be a little modified for to exclude stuff like 117 since that adds to 9

OpenStudy (anonymous):

What is so special when that adds to 9?

OpenStudy (anonymous):

I suppose all the candies will be handed out...

OpenStudy (anonymous):

n-1 C r-1 r is kids and n is candies

OpenStudy (anonymous):

@divu.mkr Can you explain it?

ganeshie8 (ganeshie8):

this is same as, finding # of ordered pairs of whole numbers x, y, z such that x+y+z = 7

ganeshie8 (ganeshie8):

u familiar wid stars and bars ?

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

"Suppose one has k objects (to be represented as stars; in the example below k = 7) to be placed into n bins (in the example n = 3), such that all bins contain at least one object;" But in this case, it is possible that the bin contains no objects.

OpenStudy (anonymous):

just remember this formula when dividing n similar things to r persons/variables

OpenStudy (anonymous):

But why it works in this case?

OpenStudy (anonymous):

because identical and similar means same!! :D

OpenStudy (anonymous):

But why n-1 and r-1 ?

OpenStudy (anonymous):

that what the formula is.... and try to memorize it rather thn derivation

OpenStudy (anonymous):

Maths is not just about remember the formula. Even though I remember it, I don't know when I can apply it, that means just memorizing is useless. Btw, the formula is for all children having at least one candy. However, it is possible that some children do not even have a candy. How can it be applied here?

OpenStudy (anonymous):

(Add the special case to the formula lol)

OpenStudy (anonymous):

if u knew that then, y did you putted up ur ques here..?

OpenStudy (anonymous):

I didn't know....

OpenStudy (anonymous):

But apart from using that stars and bars formula with adding the special case, how to solve this question. (I am not even sure if using stars and bars formula with adding the special case works.)

ganeshie8 (ganeshie8):

stars and bars works perfectly here, 7 stars, 2 bars :- ``` **|***|** ```

ganeshie8 (ganeshie8):

the \(2\) bars can be placed in available \(7+2 = 9\) slots in \(\large \binom {9}{2} = 36\) ways

ganeshie8 (ganeshie8):

So # of whole number solutions for the equation, x+y+z = 7 wud be 36

OpenStudy (anonymous):

What about ||******* ? Has it been counted?

ganeshie8 (ganeshie8):

yup bars can be anywhere

ganeshie8 (ganeshie8):

``` ||******* *******|| ```

ganeshie8 (ganeshie8):

think of 7 stars + 2 bars as slots to be filled. so 9 C 2 accounts for, ways to choose 9 slots for 2 bars

ganeshie8 (ganeshie8):

to convince ourselves, we can try below simple problem :- x+y = 7

ganeshie8 (ganeshie8):

goal is to find all whole number x, y such that , they addup to 7

ganeshie8 (ganeshie8):

we set it up using, 7 stars and 1 bar ``` ****|** ``` 8 slots, and 1 bar so # of possible solutions is 8 C 1 = 8